# Umlaut information

### v1

@@ IV.5 Operational interpretation in unassisted communication
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{\geq}}-\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)\log|\mathcal{Y}|+\mathbb{P}\left(\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}\right)\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right)\right),$
@@
  • in (i) we condition on $\mathcal{E}_{k}^{(\varepsilon)}$ and on its complementary;
  • in (ii) we make two lower bounds: we use (142) of Lemma 33 in order to lower bound the expectation value conditioned on $\mathcal{E}_{k}^{(\varepsilon)}$; then, we notice that both $v(x)\coloneqq\hat{N}^{x}_{k}$ and $u(x)\coloneqq P_{X}(x)-\varepsilon$ belong to $[0,1]^{\mathcal{X}}$ since we chose $0<\varepsilon<\min\{P_{X}(x)\,:\,x\in{\rm supp}(P_{X})\}$, and $v(x)\geq u(x)$ in $\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}$, as $\big{|}\hat{N}_{k}^{x}-P_{X}(x)\big{|}\leq\varepsilon$, hence we can use the monotonicity property of Lemma 33.
@@
$\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})\geq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right);$ (157)
@@
$\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $\geq\liminf_{\varepsilon\to 0^{+}}\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right)\right)$ (158)
@@
$U(X;Y)\geq\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})\geq U(X;Y),$ (159)
@@
Having established the achievability of the umlaut information via $E^{\emptyset}_{L}(0^{+},\pazocal{W})$ in the large list limit (Theorem 29), we can further upper bound the gap for finite list sizes $L$ as $\mathcal{O}(L^{-1/3})$.
@@
Let $\pazocal{W}$ be a discrete memoryless channel from $\mathcal{X}$ to $\mathcal{Y}$. Let $\bar{P}_{X}\in\mathcal{P}(\mathcal{X})$ be the probability distribution achieving the maximum in the definition of channel umlaut information, let $\bar{P}_{\min}\coloneqq\min_{x\in{\rm supp}\bar{P}_{X}}\bar{P}_{X}(x)$, let
$K\coloneqq\max_{y\in\mathcal{Y}_{\pazocal{W},\bar{P}_{X}}}\sum_{x\in{\rm supp}(\bar{P}_{X})}-\log\pazocal{W}(y|x)$ (160)
and let
$p^{\ast}\coloneqq\sqrt[3]{\frac{2}{L+1}\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{K}}.$ (161)
Then, we have
$0\leq U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq\begin{cases}\frac{3}{(L+1)^{1/3}}\left(\left(\log|\mathcal{Y}|+U(\pazocal{W})\right)\frac{K^{2}}{4}\right)^{1/3}&p^{\ast}<\bar{P}_{\min}\\ \frac{1}{\bar{P}_{\min}^{2}}\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{L+1}+\bar{P}_{\min}K&p^{\ast}\geq\bar{P}_{\min}\end{cases}.$ (162)
In particular, we see that, for any fixed $\pazocal{W}$, there exists $L^{\ast}>0$, defined as $L^{\ast}\coloneqq 2\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{K\bar{P}_{\min}^{3}}$, such that
$0\leq U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq\frac{3}{(L+1)^{1/3}}\left(\left(\log|\mathcal{Y}|+U(\pazocal{W})\right)\frac{K^{2}}{4}\right)^{1/3},$ (163)
when $L\geq L^{\ast}$. The upper bound vanishes for $L\to\infty$.
@@
We proceed as in the proof of Theorem 29. After fixing $P_{X}\in\mathcal{P}(\mathcal{X})$, we choose $0<\varepsilon<\min_{x\in{\rm supp}(P_{X})}P_{X}(x)$ as in the proof of Theorem 29 and we define $\mathcal{E}_{k}^{(\varepsilon)}\coloneqq\{\,\exists\,x\in\mathcal{X}\,:\,\big{|}N_{k}^{x}-P_{X}(x)\big{|}>\varepsilon\}$. We call $p_{\varepsilon}\coloneqq\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)$ to unburden the notation. Then we can go further in (156):
$\ell_{k,u_{k}}$ $(\pazocal{W},P_{X})$ (164)
$\geq-p_{\varepsilon}\log|\mathcal{Y}|+(1-p_{\varepsilon})\left(-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right)\right)$
$=-p_{\varepsilon}\log|\mathcal{Y}|+(1-p_{\varepsilon})\left(-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\left(\prod_{x\in{\rm supp}(P_{X})}\pazocal{W}(y|x)\right)^{-\varepsilon}\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)\right)$
$\geq-p_{\varepsilon}\log|\mathcal{Y}|+\varepsilon(1-p_{\varepsilon})\min_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\sum_{x\in{\rm supp}(P_{X})}\log\pazocal{W}(y|x)$
@@
$\geq-p_{\varepsilon}\log|\mathcal{Y}|+\varepsilon\min_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\sum_{x\in{\rm supp}(P_{X})}\log\pazocal{W}(y|x)+(1-p_{\varepsilon})U(X;Y),$
@@
$\max_{P_{X}}\ell_{k,u_{k}}(\pazocal{W},P_{X})-U(\bar{X};\bar{Y})\geq-p_{\varepsilon}\log|\mathcal{Y}|+\varepsilon\min_{y\in\mathcal{Y}_{\pazocal{W},\bar{P}_{X}}}\sum_{x\in{\rm supp}(\bar{P}_{X})}\log\pazocal{W}(y|x)-p_{\varepsilon}U(\bar{X};\bar{Y}),$ (165)
where $\varepsilon<\bar{P}_{\min}$, with $\bar{P}_{\min}\coloneqq\min_{x\in{\rm supp}(\bar{P}_{X})}\bar{P}_{X}(x)$. whence, setting $k=L+1$, we get
$U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq p_{\varepsilon}\left(\log|\mathcal{Y}|+U(\pazocal{W})\right)+\varepsilon K_{\pazocal{W},\bar{P}_{X}},$ (166)
@@
$K_{\pazocal{W},\bar{P}_{X}}\coloneqq\max_{y\in\mathcal{Y}_{\pazocal{W},\bar{P}_{X}}}\sum_{x\in{\rm supp}(\bar{P}_{X})}-\log\pazocal{W}(y|x).$ (167)
Now we need a quantitative upper bound to $p_{\varepsilon}$. This can be obtained by means of Chebyshev’s inequality
$\mathbb{P}\left(\big{|}N_{k}^{x}-P_{X}(x)\big{|}>\varepsilon\right)\leq\frac{{\rm Var}[N_{k}^{x}]}{\varepsilon^{2}}=\frac{{\rm Var}_{X\sim P_{X}}[\mathbbm{1}_{X=x}]}{k\varepsilon^{2}}=\frac{P_{X}(x)(1-P_{X}(x))}{k\varepsilon^{2}}$ (168)
and of the union bound
$\mathbb{P}\left(\exists\,x\in\mathcal{X}:\big{|}N_{k}^{x}-P_{X}(x)\big{|}>\varepsilon\right)\leq\sum_{x\in\mathcal{X}}\frac{P_{X}(x)(1-P_{X}(x))}{k\varepsilon^{2}}\leq\frac{1}{k\varepsilon^{2}}.$ (169)
Therefore, (166) can be upper bounded as
$U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq\frac{1}{\varepsilon^{2}}\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{L+1}+\varepsilon K_{\pazocal{W},\bar{P}_{X}}.$ (170)
For any $A,B>0$, it easy to prove that the function $f(\varepsilon)\coloneqq A/\varepsilon^{2}+B\varepsilon$, defined for $\varepsilon>0$, has an unique minimum for $\varepsilon=\varepsilon^{\ast}$, where
$\varepsilon^{\ast}=\sqrt[3]{\frac{2A}{B}}>0$ (171)
Therefore, by a simple argument about the monotonicity of $f$, we have
$\inf\{f(\varepsilon):0<\varepsilon<\varepsilon_{0}\}=\begin{cases}f(\varepsilon^{\ast})=3A^{1/3}\left(\frac{B}{2}\right)^{2/3}&\varepsilon^{\ast}<\varepsilon_{0}\\ f(\varepsilon_{0})&\varepsilon^{\ast}\geq\varepsilon_{0}\end{cases},$ (172)
whence, minimising the RHS of (170), we get
$U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq\begin{cases}\frac{3\left(\log|\mathcal{Y}|+U(\pazocal{W})\right)^{1/3}}{(L+1)^{1/3}}\left(\frac{K_{\pazocal{W},\bar{P}_{X}}}{2}\right)^{2/3}&p^{\ast}<\bar{P}_{\min}\\ \frac{1}{\bar{P}_{\min}^{2}}\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{L+1}+\bar{P}_{\min}K_{\pazocal{W},\bar{P}_{X}}&p^{\ast}\geq\bar{P}_{\min}\end{cases},$ (173)
where
$p^{\ast}\coloneqq\sqrt[3]{\frac{2}{L+1}\frac{\log|\mathcal{Y}|+U(\pazocal{W})}{K_{\pazocal{W},\bar{P}_{X}}}}$ (174)
@@
IV.6 An SDP bound on the unassisted zero-rate error exponent
@@
$E^{\emptyset}(0^{+},\pazocal{W})$ $=-\min_{P_{X}}\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}$ (175)
@@
where we have introduced $A_{xx^{\prime}}\coloneqq-\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}$, $p$ as the vector of the probability distribution $P_{X}$ and $B_{p}\coloneqq pp^{\intercal}$. $B$ belongs to the set of completely positive matrices $\mathrm{CP}_{|\mathcal{X}|}$, defined as
$\mathrm{CP}_{d}\coloneqq\left\{B\in\mathbb{R}^{d\times d}\,:\,B=\sum_{i=1}^{k}v_{i}v_{i}^{\intercal},\,v_{i}\in\mathbb{R}^{d}_{+},\,k\in\mathbb{N}\right\},$ (176)
where +$\mathbb{R}^{d}_{+}$ is the cone of vectors in $\mathbb{R}^{d}$ with non-negative entries. Let $u_{d}$ be the uniform vector in $\mathbb{R}^{d}$ given by $u_{d}=(1,\dots,1)^{\intercal}$. It holds that $u_{|\mathcal{X}|}^{\intercal}B_{p}u_{|\mathcal{X}|}=1$. Conversely, let us suppose that $B\in cp_{|\mathcal{X}|}$ and $u_{|\mathcal{X}|}^{\intercal}Bu_{|\mathcal{X}|}=1$. To unburden the notation, from now on the dimension of vectors and matrices will be implied. On one hand, for some $k\in\mathbb{N}$,
$1=u^{\intercal}Bu=\sum_{i=1}^{k}(u^{\intercal}v_{i})^{2}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\sum_{i=1}^{k}q_{i},$ (177)
@@
$B=\sum_{i=1}^{k}q_{i}\frac{v_{i}}{u^{\intercal}v_{i}}\frac{v_{i}^{\intercal}}{u^{\intercal}v_{i}}=\sum_{i=1}^{k}q_{i}p_{i}p_{i}^{\intercal},$ (178)
@@
$\operatorname{Tr}[BA]=\sum_{i=1}^{k}q_{i}p_{i}^{\intercal}Ap_{i}\leq\max_{p}p^{\intercal}Ap=\max_{p}\operatorname{Tr}[B_{p}A].$ (179)
@@
$E^{\emptyset}(0^{+},\pazocal{W})=\max_{\begin{subarray}{c}B\in\mathrm{CP}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[AB].$ (180)
@@
$\mathrm{CP}_{d}\subseteq\mathrm{NN}_{d}\cap\mathrm{PSD}_{d},$ (181)
@@
$\mathrm{CP}_{d}=\mathrm{NN}_{d}\cap\mathrm{PSD}_{d},$ (182)
if and only if $d\leq 4$ [32, Theorem 2.4, Example 2.7]. We have therefore proved the following statement.
@@
$E^{\emptyset}(0^{+},\pazocal{W})=\max_{\begin{subarray}{c}B\in\mathrm{CP}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[A^{(\pazocal{W})}B],$ (183)
@@
$A_{xx^{\prime}}^{(\pazocal{W})}\coloneqq-\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}.$ (184)
@@
$E^{\emptyset}(0^{+},\pazocal{W})\leq\max_{\begin{subarray}{c}B\in\mathrm{PSD}\cap\mathrm{NN}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[A^{(\pazocal{W})}B]$ (185)

### v2

@@ IV.6 Operational interpretation in unassisted communication
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{\geq}}\mathbb{P}\left(\mathcal{E}_{k}^{(\varepsilon)}\right)\cdot 0+\mathbb{P}\left(\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}\right)\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right)\right),$
@@
  • in (i) we condition on $\mathcal{E}_{k}^{(\varepsilon)}$ and on its complement;
  • in (ii) we introduce two lower bounds: since $v(x)\coloneqq\hat{N}^{x}_{k}$ is a probability distribution on $\mathcal{X}$ (i.e. $\|v\|_{1}=1$) we use (184) of Lemma 38 in order to lower bound the expectation value conditioned on $\mathcal{E}_{k}^{(\varepsilon)}$; then, we notice that not only $v(x)$, but also $u(x)\coloneqq P_{X}(x)-\varepsilon$ belongs to $0,1]^{\mathcal{X}}$ since we chose $0<\varepsilon<\min\{P_{X}(x)\,:\,x\in{\rm supp}(P_{X})\}$, and $v(x)\geq u(x)$ in $\big{(}\mathcal{E}_{k}^{(\varepsilon)}\big{)}^{c}$, as $\big{|}\hat{N}_{k}^{x}-P_{X}(x)\big{|}\leq\varepsilon$, hence we can use the monotonicity property of Lemma 38.
@@
$\hskip 0.0pt\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})\geq-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right);$ (199)
@@
$\hskip 0.0pt\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $\geq\liminf_{\varepsilon\to 0^{+}}\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}(P_{X}(x)-\varepsilon)\log\pazocal{W}(y|x)\right)\right)$ (200)
@@
$\hskip 0.0ptU(X;Y)\geq\liminf_{k\to\infty}\ell_{k,u_{k}}(\pazocal{W},P_{X})\geq U(X;Y),$ (201)
@@
Having established the achievability of the umlaut information via $E^{\emptyset}_{L}(0^{+},\pazocal{W})$ in the large list limit (Theorem 34), we can further upper bound the gap for finite list sizes $L$ as $O\big{(}\sqrt{\log L/L}\big{)}$.
@@
Let $\pazocal{W}$ be a discrete memoryless channel from $\mathcal{X}$ to $\mathcal{Y}$. Let $\bar{P}_{X}\in\mathcal{P}(\mathcal{X})$ be the probability distribution achieving the maximum in the definition of channel umlaut information, set $\bar{p}_{\min}\coloneqq\min_{x\in{\rm supp}\left(\bar{P}_{X}\right)}\bar{P}_{X}(x)$, define $\pazocal{W}_{\min}\coloneqq\min\{\pazocal{W}(y|x):x\in\mathcal{X},\,y\in\mathcal{Y},\,\pazocal{W}(y|x)\neq 0\}$ and pick some $\varepsilon\in(0,1)$. Then, we have
$\hskip 0.0pt0\leq U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq|\mathcal{X}|\exp\left(-\frac{\varepsilon^{2}}{2}\bar{p}_{\min}(L+1)\right)-\varepsilon\log\pazocal{W}_{\min}.$ (202)
In particular, calling
$\hskip 0.0pt\varepsilon(L)=\sqrt{\frac{1}{\bar{p}_{\min}}\frac{\log(L+1)}{L+1}},$ (203)
we can set $\varepsilon=\min\{\varepsilon(L),1/2\}$, getting
$\hskip 0.0pt|U(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})|=O\left(\left(\frac{\log L}{L}\right)^{1/2}\right)$ (204)
for $L\to\infty$.
@@
We proceed similarly to the proof of Theorem 34. After fixing $P_{X}\in\mathcal{P}(\mathcal{X})$, let $\varepsilon\in(0,1)$ and, for every $x\in\mathrm{supp}(P_{X}),$ let $\mathcal{F}_{k,x}^{(\varepsilon)}$ be the event $\hat{N}_{k}^{x}\leq P_{X}(x)(1-\varepsilon)$. Since $k\hat{N}^{k}_{x}$, by the very definition of $\hat{N}_{k}^{x}$, has a binomial distribution (see eq. (195)), by the Chernoff bound on the lower tail, we have
$\hskip 0.0pt\mathbb{P}\left(\mathcal{F}_{k,x}^{(\varepsilon)}\right)=\mathbb{P}\left(\hat{N}_{k}^{x}\leq P_{X}(x)(1-\varepsilon)\right)\leq\exp\left(-\frac{\varepsilon^{2}}{2}P_{X}(x)k\right).$ (205)
Therefore, the event
$\hskip 0.0pt\mathcal{F}_{k}^{(\varepsilon)}=\left\{\exists x\in\mathrm{supp}(P_{X})\,:\,\hat{N}_{k}^{x}\leq P_{X}(x)(1-\varepsilon)\right\}=\bigcup_{x\in\mathrm{supp}(P_{X})}\mathcal{F}_{k,x}^{(\varepsilon)}$ (206)
has probability at most
$\hskip 0.0pt\mathbb{P}\left(\mathcal{F}_{k}^{(\varepsilon)}\right)\leq\sum_{x\in\mathrm{supp}(P_{X})}\exp\left(-\frac{\varepsilon^{2}}{2}P_{X}(x)k\right)\leq|\mathcal{X}|\exp\left(-\frac{\varepsilon^{2}}{2}p_{\min}k\right),$ (207)
where $p_{\min}=\min_{x\in{\rm supp}(P_{X})}P_{X}(x)$. We call $p_{\varepsilon}\coloneqq\mathbb{P}\left(\mathcal{F}_{k}^{(\varepsilon)}\right)$ to unburden the notation. Similarly to (198):
$\hskip 0.0pt\ell_{k,u_{k}}(\pazocal{W},P_{X})$ $=\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\right]$ (208)
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\,\middle|\,\mathcal{E}_{k}^{(\varepsilon)}\,\right]p_{\varepsilon}$
$\quad+\mathbb{E}_{X^{k}\sim P_{X}^{k}}\left[-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}\hat{N}_{k}^{x}\log\pazocal{W}(y|x)\right)\,\middle|\,(\mathcal{E}_{k}^{(\varepsilon)})^{c}\,\right](1-p_{\varepsilon})$
$\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(ii)}}}}{{\geq}}\left(-\log\sum_{y\in\mathcal{Y}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)(1-\varepsilon)\log\pazocal{W}(y|x)\right)\right)(1-p_{\varepsilon}),$
where
  • in (i) we condition on $\mathcal{F}_{k}^{(\varepsilon)}$ and on its complement;
  • in (ii) we make two lower bounds: since $v(x)\coloneqq\hat{N}^{x}_{k}$ is a probability distribution on $\mathcal{X}$ (i.e. $\|v\|_{1}=1$) we use (184) of Lemma 38 in order to lower bound with zero the expectation value conditioned on $\mathcal{F}_{k}^{(\varepsilon)}$; then, we notice that $0,1]^{\mathcal{X}}$ $v(x)\geq u(x)\coloneqq P_{X}(x)(1-\varepsilon)$ in $\big{(}\mathcal{F}_{k}^{(\varepsilon)}\big{)}^{c}$, hence we can use the monotonicity property of Lemma 38.
Now, we can go further:
$\hskip 0.0pt\ell_{k,u_{k}}$ $(\pazocal{W},P_{X})$ (209)
$\geq(1-p_{\varepsilon})\left(-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\exp\left(\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)(1-\varepsilon)\log\pazocal{W}(y|x)\right)\right)$
$=(1-p_{\varepsilon})\left(-\log\sum_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\exp\left(-\varepsilon\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)\log\pazocal{W}(y|x)\right)\exp\left(\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)\right)\right)$
$\geq\varepsilon(1-p_{\varepsilon})\min_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\sum_{x\in{\rm supp}(P_{X})}P_{X}(x)\log\pazocal{W}(y|x)$
@@
$\geq\varepsilon\min_{y\in\mathcal{Y}_{\pazocal{W},P_{X}}}\sum_{x\in\mathcal{X}}P_{X}(x)\log\pazocal{W}(y|x)+(1-p_{\varepsilon})U(X;Y),$
@@
$\hskip 0.0pt\max_{P_{X}}\ell_{k,u_{k}}(\pazocal{W},P_{X})-U(\bar{X};\bar{Y})\geq\varepsilon\min_{y\in\mathcal{Y}_{\pazocal{W},\bar{P}_{X}}}\sum_{x\in{\rm supp}(\bar{P}_{X})}\bar{P}_{X}(x)\log\pazocal{W}(y|x)-p_{\varepsilon}U(\bar{X};\bar{Y}),$ (210)
whence, setting $k=L+1$, we get
$\hskip 0.0ptU(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq p_{\varepsilon}U(\pazocal{W})+\varepsilon K_{\pazocal{W},\bar{P}_{X}},$ (211)
@@
$\hskip 0.0ptK_{\pazocal{W},\bar{P}_{X}}$ $\coloneqq\max_{y\in\mathcal{Y}_{\pazocal{W},\bar{P}_{X}}}\sum_{x\in{\rm supp}(\bar{P}_{X})}-\bar{P}_{X}(x)\log\pazocal{W}(y|x)\leq-\log\pazocal{W}_{\min},$ (212)
with $\pazocal{W}_{\min}\coloneqq\min\{\pazocal{W}(y|x):x\in\mathcal{X},\,y\in\mathcal{Y},\,\pazocal{W}(y|x)\neq 0\}$. Using the estimate on $p_{\varepsilon}$ and calling $\bar{p}_{\min}\coloneqq\min_{x\in{\rm supp}(\bar{P}_{X})}\bar{P}_{X}(x)$, (211) can be upper bounded as
$\hskip 0.0ptU(\pazocal{W})-E^{\emptyset}_{L}(0^{+},\pazocal{W})\leq|\mathcal{X}|\exp\left(-\frac{\varepsilon^{2}}{2}\bar{p}_{\min}(L+1)\right)-\varepsilon\log\pazocal{W}_{\min},$ (213)
@@
IV.7 An SDP bound on the unassisted zero-rate error exponent
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})$ $=-\min_{P_{X}}\sum_{x,x^{\prime}\in\mathcal{X}}P_{X}(x)P_{X}(x^{\prime})\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}$ (214)
@@
where we have introduced $A_{xx^{\prime}}\coloneqq-\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}$, $p$ as the vector of the probability distribution $P_{X}$ and $_{p}\coloneqq pp^{\intercal}. BB$ belongs to the set of completely positive matrices $\mathrm{CP}_{|\mathcal{X}|}$, defined as
$\hskip 0.0pt\mathrm{CP}_{d}\coloneqq\left\{B\in\mathbb{R}^{d\times d}\,:\,B=\sum_{i=1}^{k}v_{i}v_{i}^{\intercal},\,v_{i}\in\mathbb{R}^{d}_{+},\,k\in\mathbb{N}\right\},$ (215)
where $\mathbb{R}^{d}_{+}$ is the cone of vectors in $\mathbb{R}^{d}$ with non-negative entries. Let $u_{d}$ be the uniform vector in $\mathbb{R}^{d}$ given by $u_{d}=(1,\dots,1)^{\intercal}$. It holds that $u_{|\mathcal{X}|}^{\intercal}B_{p}u_{|\mathcal{X}|}=1$. Conversely, let us suppose that $B\in cp_{|\mathcal{X}|}$ and $u_{|\mathcal{X}|}^{\intercal}Bu_{|\mathcal{X}|}=1$. To unburden the notation, from now on the dimension of vectors and matrices will be implied. On the one hand, for some $k\in\mathbb{N}$,
$\hskip 0.0pt1=u^{\intercal}Bu=\sum_{i=1}^{k}(u^{\intercal}v_{i})^{2}\stackrel{{\scriptstyle\mathclap{\mbox{\scriptsize(i)}}}}{{=}}\sum_{i=1}^{k}q_{i},$ (216)
@@
$\hskip 0.0ptB=\sum_{i=1}^{k}q_{i}\frac{v_{i}}{u^{\intercal}v_{i}}\frac{v_{i}^{\intercal}}{u^{\intercal}v_{i}}=\sum_{i=1}^{k}q_{i}p_{i}p_{i}^{\intercal},$ (217)
@@
$\hskip 0.0pt\operatorname{Tr}[BA]=\sum_{i=1}^{k}q_{i}p_{i}^{\intercal}Ap_{i}\leq\max_{p}p^{\intercal}Ap=\max_{p}\operatorname{Tr}[B_{p}A].$ (218)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})=\max_{\begin{subarray}{c}B\in\mathrm{CP}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[AB].$ (219)
@@
$\hskip 0.0pt\mathrm{CP}_{d}\subseteq\mathrm{NN}_{d}\cap\mathrm{PSD}_{d},$ (220)
@@
$\hskip 0.0pt\mathrm{CP}_{d}=\mathrm{NN}_{d}\cap\mathrm{PSD}_{d},$ (221)
if and only if $d\leq 4$ [35, Theorem 2.4, Example 2.7]. We have therefore proved the following statement.
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})=\max_{\begin{subarray}{c}B\in\mathrm{CP}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[A^{(\pazocal{W})}B],$ (222)
@@
$\hskip 0.0ptA_{xx^{\prime}}^{(\pazocal{W})}\coloneqq-\log\sum_{y\in\mathcal{Y}}\sqrt{\pazocal{W}(y|x)\pazocal{W}(y|x^{\prime})}.$ (223)
@@
$\hskip 0.0ptE^{\emptyset}(0^{+},\pazocal{W})\leq\max_{\begin{subarray}{c}B\in\mathrm{PSD}\cap\mathrm{NN}\\ u^{\intercal}Bu=1\end{subarray}}\operatorname{Tr}[A^{(\pazocal{W})}B]$ (224)